Since astronauts in orbit are apparently weightien, a clever method of measuring their masses is needed to monitor their man gains or losses to adjust diets. One way to do this is to exert a known force on an astronauts and measure the acceleration produced. Suppose a net external force of 50.0 N is exerted and the astronaut's acceleration measured to be 0.893 m/s², (a) calculate her mass: (b) by exerting a force on the astronaut, the vehicle in which they orbit experiences an equal and opposite force. Discuss how this would affect the measurement of the astronaut's acceleration. Proposed a method in which recoil of the vehicle is avoided.


1
Expert's answer
2021-10-06T11:45:41-0400

(a) Mass:


"a=\frac {F_\text{net}}{m}=56\text{ kg}."

(b) Yes, the magnitude of the measured acceleration will reduce. A method in which recoil of the vehicle is avoided may take conservation of momentum into account: the astronaut holds a gun that ejects a projectile of known momentum. By measuring the initial velocity of the astronaut, it is possible to determine the mass:


"0=m_pv_p-m_av_a,\\\space\\
m_a=m_p\frac{v_p}{v_a}."